Represent (2+ root5 )on the real line.
1
Expert's answer
2020-03-09T13:55:28-0400

We start by drawing equal spaced partitions on the number line to represent integers.

Using Pythagoras theorem, we know that "\sqrt 5" is the length of hypotenuse of the triangle whose other two sides are , 1 units and 2 units. Thus we create a right angled triangle with its acute angled vertex lying at 2 on the number line. Finally, on construction of the right triangle, we use the compass and cut an arc of length "\sqrt 5" keeping the fixed end of the compass on 2 on the number line. The point where the arc cuts the number line is "2 + \sqrt 5" .